OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 7. Find the number(s).

OpenStudy (anonymous):

x+(6/x) = 7 I think that is what you are saying then you solve for x

OpenStudy (anonymous):

i assume?

OpenStudy (anonymous):

6 = 7x-x^2 x^2-7x+6 = 0 (x-1)(x-6) so your answers are 1 and 6 I think

OpenStudy (anonymous):

ohh okay thanks

OpenStudy (anonymous):

Do you see what I did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 7. Let the number = x Its reciprocal = 1/x 1 So x + 6*---- = 7 x Now multiply throughout by x x^2 + 6 = 7x gives x^2 - 7x + 6 = 0 solve this quadratic equation

OpenStudy (anonymous):

okay 1 sec

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

got 6,1

OpenStudy (anonymous):

correct!!

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

haha for real?

OpenStudy (anonymous):

lol what?

OpenStudy (anonymous):

People don't read before they post, and you get 2 or 3 people posting the exact same answer to a question.

OpenStudy (anonymous):

lol ohh yea i see